# Information Density Bounds for Privacy

### v1

@@ III-D Relationship to Maximal Guesswork Leakage
A central example of a cost function is the expected number of attempts (or moments thereof) required to correctly guess a secret [49, 50]. In this context, a guessing function $G$ is a one-to-one mapping from $\mathcal{X}$ to $[\lvert\mathcal{X}\rvert]$, and the guesswork is the minimum expected number of attempts required to correctly identify $X$, i.e., $\min_{G}\mathbb{E}[G(X)]$. Recently, guesswork has been used to define two information measures called maximal guesswork leakage and pointwise maximal guesswork leakage [23]. Both measures describe adversaries aiming to minimize the guesswork associated with randomized functions of $X$ and compare the prior guesswork with the posterior guesswork. Per usual terminology, maximal guesswork leakage describes the average amount of information leaking through a mechanism, while pointwise maximal guesswork leakage is defined for each outcome $y\in\mathcal{Y}$.
@@
Definition 10 (Pointwise maximal guesswork leakage [23, Def. 6]).
@@
Pointwise maximal guesswork leakage is a special case of the cost function model of PMC. In particular, given a random variable $U$ satisfying the Markov chain $U-X-Y$ and a guessing function $G:\mathcal{U}\to[\lvert\mathcal{U}\rvert]$, there exists a cost function $d:\mathcal{U}\times\mathcal{W}\to[\lvert\mathcal{U}\rvert]$ such that
@@
(see [23, Sec. III] for the exact construction). Furthermore, we showed in (the proof of) Proposition 2 that for each $U$ satisfying the Markov chain $U-X-Y$ and each cost function $d:\mathcal{U}\times\mathcal{W}\to\mathbb{R}_{+}$, there exists a cost function $c:\mathcal{X}\times\mathcal{W}\to\mathbb{R}_{+}$ such that
@@
$\mathcal{L}^{\mathsf{G-pw}}(X\to y)=\Lambda(X\to y)=D_{\infty}(P_{X}\|P_{X\mid Y=y}),\quad y\in\mathcal{Y}.$
@@ III-E PMC on General Alphabets
Let $\mathcal{X}$ be a finite set with cardinality $k$, and let $\Delta^{k-1}$ denote the set of all distributions on $\mathcal{X}$, i.e., the $(k-1)$-dimensional probability simplex. Liao et al. [31] introduced a class of loss functions called $\alpha$-loss with $\alpha\in[1,\infty]$ that interpolate between the log-loss at $\alpha=1$ and the 0-1 loss at $\alpha=\infty$. More precisely, they defined the $\alpha$-loss $c_{\alpha}:\mathcal{X}\times\Delta^{k-1}\to\bar{}\mathbb{R}_{+}$ by
@@
where $Q\in\Delta^{k-1}$ is a distribution on $\mathcal{X}$. Then, they used the $\alpha$-loss to define a class of tunable information leakage measure called (maximal) $\alpha$-leakage [31, Def. 5 and 6]. [^2] Note that (maximal) $\alpha$-leakage assumes that the adversary maximizes an objective function such that the optimal action corresponds to minimizing the $\alpha$-loss. In contrast, here we assume that the adversary directly minimizes the $\alpha$-loss.
@@ IV-A Properties of Pointwise Maximal Cost
$\Lambda_{P_{XY\mid Z}}(X\to y\mid z)\coloneqq\log\sup_{\begin{subarray}{c}(\mathcal{W},\mathcal{S}_{\mathcal{W}}),\\ c\in\mathcal{C}_{z}\end{subarray}}\;\frac{\inf_{P_{W\mid Z}}\;\mathbb{E}\Big{[}c(X,W)\mid Z=z\Big{]}}{\inf_{P_{\hat{W}\mid Y,Z}}\;\mathbb{E}\Big{[}c(X,\hat{W})\mid Y=y,Z=z\Big{]}},$ (27)
@@
As a special case, if the Markov chain $Z-X-Y$ holds, then $\Lambda_{P_{XY\mid Z}}(X\to y\mid z)=D_{\infty}(P_{X\mid Z=z}\|P_{X\mid Y=y})$. This scenario corresponds to the important case where the side information determines the adversary’s prior knowledge about $X$ but does not influence the mechanism through which information is released.
@@
  5. 5.
    (Pre-processing). Suppose the Markov chain $Z-X-Y$ holds. Then, for all $y\in\mathcal{Y}$ it holds that
@@
  6. 6.
    (Post-processing). Suppose the Markov chain $X-Y-Z$ holds. Then, we have
@@
    $\Lambda_{P_{XY_{1}Y_{2}}}(X\to y_{1},y_{2})\leq\Lambda_{P_{XY_{1}}}(X\to y_{1})+\Lambda_{P_{XY_{2}\mid Y_{1}}}(X\to y_{2}\mid y_{1}),$
@@ IV-B Examples
$\varepsilon=\max_{j\in[n]}\Lambda(X\to j)=\log\;\Big{(}1+\max_{j\in[n]}P_{X}(j)(e^{\varepsilon_{r}}-1)\Big{)}.$ (28)
Note that $\varepsilon\leq\varepsilon_{r}$ for all distributions $P_{X}$. Interestingly, $\varepsilon\to\varepsilon_{r}$ as $\max_{j\in[n]}P_{X}(j)\to 1$.
@@
Suppose $X$ is a finite random variable with alphabet $\mathcal{X}=[n]$. Recall from Section II-B that PML-extremal mechanisms constitute a class of optimal mechanisms maximizing sub-convex utility functions [62, Def. 5]. Let $0\leq\varepsilon_{u}<\log\,\frac{1}{1-p_{\mathrm{min}}}$. This range of values for $\varepsilon_{u}$ corresponds to the high-privacy regime. Extremal mechanisms in the high-privacy regime can be expressed as
@@
$\Lambda(X\to j)=\log\;\max_{i\in[n]}\;\frac{P_{Y}(j)}{P_{Y\mid X=i}(j)}=\log\,\frac{P_{X}(j)}{1-e^{\varepsilon_{u}}\Big{(}1-P_{X}(j)\Big{)}}.$
@@
Below, we evaluate the PMC from each data point to the released outcome. Due to the symmetry of the problem, the Laplace mechanism leaks the same amount of information about all data points, and for notational convenience, we calculate $\Lambda(X_{n}\to y)$ with $y\in\mathbb{R}$. Our analysis in this example complements a similar analysis performed with the Laplace mechanism and PML in [21, Prop. 4.5].
@@
First, we argue that to find the largest PMC across the $y$’s, it suffices to consider $y\geq d$ and $y\leq c$. To see why, note that in the numerator, using the fact that $-\lvert a\rvert\leq\min\{-a,a\}$ for all $a\in\mathbb{R}$ we have
$\mathbb{E}_{X^{n}}\left[\exp\left(-\frac{\lvert y-\frac{\sum_{i=1}^{n}X_{i}}{n}\rvert}{b}\right)\right]\leq\min\left\{\mathbb{E}_{X^{n}}\left[\exp\left(-\frac{y-\frac{\sum_{i=1}^{n}X_{i}}{n}}{b}\right)\right],\mathbb{E}_{X^{n}}\left[\exp\left(\frac{y-\frac{\sum_{i=1}^{n}X_{i}}{n}}{b}\right)\right]\right\},$
for all $y\in\mathbb{R}$. Moreover, in the denominator, the mapping $y\mapsto\mathbb{E}_{X^{n-1}}\left[\exp\left(-\frac{\lvert y-\frac{x_{n}}{n}-\frac{\sum_{i=1}^{n-1}X_{i}}{n}\rvert}{b}\right)\right]$ is increasing in $(-\infty,c]$ and decreasing in $[d,\infty)$ for all $x_{n}\in[c,d]$.
@@
$\mathbb{E}_{X^{n}}\left[\exp\left(-\frac{y-\frac{\sum_{i=1}^{n}X_{i}}{n}}{b}\right)\right]$ $=\exp\left(-\frac{y}{b}\right)\;\mathbb{E}_{X^{n}}\left[\exp\left(\frac{\sum_{i=1}^{n}X_{i}}{nb}\right)\right]$
@@
$\mathbb{E}_{X^{n-1}}\left[\exp\left(-\frac{\displaystyle y-\frac{x_{n}}{n}-\frac{\sum_{i=1}^{n-1}X_{i}}{n}}{b}\right)\right]$ $=\exp\left(-\frac{y}{b}\right)\cdot\exp\left(\frac{x_{n}}{nb}\right)\;\mathbb{E}_{X^{n-1}}\left[\exp\left(\frac{\sum_{i=1}^{n-1}X_{i}}{nb}\right)\right]$
@@
[…]
@@
$\log\;\mathbb{E}_{X}\left[\exp\left(\frac{2y(A+X)}{2\sigma^{2}}\right)\right]\leq\Lambda(X\to y)\leq\log\;\mathbb{E}_{X}\left[\exp\left(\frac{2y(A+X)+A^{2}}{2\sigma^{2}}\right)\right],$
@@
$-\frac{Ay}{\sigma^{2}}\leq\Lambda(X\to y)\leq\frac{A(A-4y)}{2\sigma^{2}}.$ (39)
@@
$\mathbb{P}\Big{\{}\Lambda(X\to Y)\geq\beta+\frac{A^{2}}{2\sigma^{2}}\Big{\}}$$\leq\mathbb{P}\left\{\frac{A(A+4\lvert Y\rvert)}{2\sigma^{2}}\geq\beta+\frac{A^{2}}{2\sigma^{2}}\right\}$ (41)
=$=\mathbb{P}\left\{\lvert Y\rvert\geq\frac{\sigma^{2}\beta}{2A}\right\}$ (42)
@@ V PMC, PML, and Connections to Other Privacy Measures
In general, the upper and lower bounds on the information density are not independent quantities. These bounds influence each other so that imposing a lower bound on the information density may imply a certain upper bound, and vice versa. In the context of privacy, this naturally raises the question of whether a PMC guarantee inherently provides a certain PML guarantee, or whether a PML guarantee implies a certain PMC guarantee.
In this section, we explore how PMC and PML guarantees interact with each other and with other privacy definitions, namely, ALIP, LIP, and LDP. First, in Section V-A, we assume that $X$ is a finite random variable and discuss how various privacy guarantees interrelate. We also highlight the implications of our results for designing mechanisms with optimal utility. Then, in Section V-B, we extend our discussions to secrets on standard Borel spaces.
@@ V-A Secrets with Finite Alphabets
Suppose $X\sim P_{X}$ is a finite random variable, and recall that $p_{\mathrm{min}}=\min_{x\in\mathcal{X}}P_{X}(x)$. We begin by illustrating how an LDP privacy guarantee translates into a PMC guarantee. This result combined with Proposition 1 paints a complete picture depicting how LDP restricts the information density. Proposition 3 is proved in Appendix E.
@@
Theorem 4 is proved in Appendix F.
@@
The condition $\varepsilon_{u}<\log\frac{1}{1-p_{\min}}$ restricts us to the high-privacy regime of PML and is necessary for $\varepsilon_{l}^{*}$ to be well-defined. This is because a mechanism $P_{Y\mid X}$ satisfying $\varepsilon_{u}$-PML with $\varepsilon_{u}\geq\log\frac{1}{1-p_{\min}}$ may assign $P_{Y\mid X=x}(y)=0$ for some $(x,y)$ pair with positive probability. Such a mechanism does not satisfy $\varepsilon_{l}$-PMC for any finite value of $\varepsilon_{l}$. This is also why an $\varepsilon_{u}$-PML guarantee yields useful LIP, ALIP and LDP guarantees only if $\varepsilon_{u}<\log\frac{1}{1-p_{\min}}$ (see Corollary 3).
@@
Figure 1 illustrates the bounds presented in Theorem 4. Consider a pair $(\varepsilon_{u},\varepsilon_{l})\geq(0,0)$ with $\varepsilon_{u}<\log\frac{1}{1-p_{\mathrm{min}}}$. Theorem 4 states that if $\varepsilon_{l}>\varepsilon_{l}^{*}$, then $\varepsilon_{u}$-PML is a strictly stronger privacy guarantee compared to $\varepsilon_{l}$-PMC. Such pairs constitute the gray area above the blue curve in Figure 1(a). Similarly, if $\varepsilon_{u}>\varepsilon_{u}^{*}$ then $\varepsilon_{l}$-PMC is a strictly stronger privacy guarantee compared to $\varepsilon_{u}$-PML. These pairs correspond to the gray area below the orange curve in Figure 1(a). In contrast, the blue region in Figure 1(a) represents $(\varepsilon_{u},\varepsilon_{l})$ pairs that satisfy $\varepsilon_{u}^{*}\geq\varepsilon_{u}$ and $\varepsilon_{l}^{*}\geq\varepsilon_{l}$. In this region, neither the $\varepsilon_{u}$-PML nor the $\varepsilon_{l}$-PMC privacy guarantee is stronger than the other. Interestingly, when $X$ is an equiprobable Bernoulli random variable, the gap between the $\varepsilon_{u}$-PML and $\varepsilon_{l}$-PMC curves (i.e., the blue region) disappears, as illustrated in Figure 1(b). We formalize this observation in Corollary 4.
@@
Let us now discuss some of the implications of Theorem 4. The primary consequence of this theorem is that both PML and PMC privacy guarantees translate into ALIP, LIP, and LDP guarantees. The ALIP and LIP guarantees simply follow from their definitions and the LDP guarantees follow from [19, Prop. 1], where it is shown that $(\varepsilon_{l},\varepsilon_{u})$-ALIP implies $(\varepsilon_{l}+\varepsilon_{u})$-LDP.
@@
[…]
  2. 2.
@@
[…]

### v2

@@ III-D Relationship to Maximal Guesswork Leakage
A central example of a cost function is the expected number of attempts (or moments thereof) required to correctly guess a secret [46, 45]. In this context, a guessing function $G$ is a one-to-one mapping from $\mathcal{X}$ to $\lvert\mathcal{X}\rvert$], and the guesswork is the minimum expected number of attempts required to correctly identify $X$, i.e., $\min_{G}\mathbb{E}[G(X)]$. Recently, guesswork has been used to define two information measures called maximal guesswork leakage and pointwise maximal guesswork leakage [36]. Both measures describe adversaries aiming to minimize the guesswork associated with randomized functions of $X$ and compare the prior guesswork with the posterior guesswork. Per usual terminology, maximal guesswork leakage describes the average amount of information leaking through a mechanism, while pointwise maximal guesswork leakage is defined for each outcome $y\in\mathcal{Y}$.
@@
Definition 10 (Pointwise maximal guesswork leakage [36, Def. 6]).
@@
Pointwise maximal guesswork leakage is a special case of the cost function model of PMC. In particular, given a random variable $U$ satisfying the Markov chain UXYU-X-Y and a guessing function $G:\mathcal{U}\to[\lvert\mathcal{U}\rvert$], there exists a cost function $d:\mathcal{U}\times\mathcal{W}\to[\lvert\mathcal{U}\rvert$] such that
@@
(see [36, Sec. III] for the exact construction). Furthermore, we showed in (the proof of) Proposition 2 that for each $U$ satisfying the Markov chain UXYU-X-Y and each cost function $d:\mathcal{U}\times\mathcal{W}\to\mathbb{R}_{+}$, there exists a cost function $c:\mathcal{X}\times\mathcal{W}\to\mathbb{R}_{+}$ such that
@@
$\mathcal{L}^{\mathsf{G-pw}}(X\to y)=\Lambda(X\to y)=D_{\infty}(P_{X}\|P_{X\mid Y=y}),\;y\in\mathcal{Y}.$
@@ III-E PMC on General Alphabets
Let $\mathcal{X}$ be a finite set with cardinality $k$, and let $\Delta^{k-1}$ denote the set of all distributions on $\mathcal{X}$, i.e., the (k1)(k-1)-dimensional probability simplex. Liao et al. [39] introduced a class of loss functions called $\alpha$-loss with $\alpha\in[1,\infty]$ that interpolate between the log-loss at $\alpha=1$ and the 0-1 loss at $\alpha=\infty$. More precisely, they defined the $\alpha$-loss $c_{\alpha}:\mathcal{X}\times\Delta^{k-1}\to\bar{}\mathbb{R}_{+}$ by
@@
where $Q\in\Delta^{k-1}$ is a distribution on $\mathcal{X}$. Then, they used the $\alpha$-loss to define a class of tunable information leakage measure called (maximal) $\alpha$-leakage [39, Def. 5 and 6]. [^1] Note that (maximal) $\alpha$-leakage assumes that the adversary maximizes an objective function such that the optimal action corresponds to minimizing the $\alpha$-loss. In contrast, here we assume that the adversary directly minimizes the $\alpha$-loss.
@@ IV-A Properties of Pointwise Maximal Cost
$\Lambda_{P_{XY\mid Z}}(X\to y\mid z)$ (32)
$\coloneqq\log\sup_{\begin{subarray}{c}(\mathcal{W},\mathcal{S}_{\mathcal{W}}),\\ c\in\mathcal{C}_{z}\end{subarray}}\;\frac{\inf_{P_{W\mid Z}}\;\mathbb{E}\Big[c(X,W)\mid Z=z\Big]}{\inf_{P_{\hat{W}\mid Y,Z}}\mathbb{E}\Big[c(X,\hat{W})\mid Y=y,Z=z\Big]},$ (33)
@@
As a special case, if the Markov chain ZXYZ-X-Y holds, then $\Lambda_{P_{XY\mid Z}}(X\to y\mid z)=D_{\infty}(P_{X\mid Z=z}\|P_{X\mid Y=y})$. This scenario corresponds to the important case where the side information determines the adversary’s prior knowledge about $X$ but does not influence the mechanism through which information is released.
@@
  5. (v)
    (Pre-processing). Suppose the Markov chain ZXYZ-X-Y holds. Then, for all $y\in\mathcal{Y}$ it holds that
@@
  6. (vi)
    (Post-processing). Suppose the Markov chain XYZX-Y-Z holds. Then, we have
@@
    $\Lambda_{P_{XY_{1}Y_{2}}}(X\to y_{1},y_{2})\\ \leq\Lambda_{P_{XY_{1}}}(X\to y_{1})+\Lambda_{P_{XY_{2}\mid Y_{1}}}(X\to y_{2}\mid y_{1}),$
@@ IV-B Examples
$\varepsilon=\max_{j\in[n]}\Lambda(X\to j)=\log\;\Big(1+p_{\max}\,(e^{\varepsilon_{r}}-1)\Big),$ (34)
where $p_{\max}\coloneqq\max_{j\in[n]}P_{X}(j)$. Note that $\varepsilon\leq\varepsilon_{r}$ for all distributions $P_{X}$, and $\varepsilon\to\varepsilon_{r}$ as $p_{\max}\to 1$.
Next, we compare the utility of the generalized randomized response mechanism when its parameter is tuned to satisfy a target $\varepsilon$-PMC, $\varepsilon$-PML, or $\varepsilon$-LDP privacy guarantee. In [25], we showed that the randomized response mechanism with parameter $\varepsilon_{r}$ satisfies $\log\left(\frac{e^{\varepsilon_{r}}}{p_{\min}\,(e^{\varepsilon_{r}}-1)+1}\right)$-PML, where $p_{\min}=\min_{j\in[n]}P_{X}(j)$. Recall that PML is always upper bounded by $\log\frac{1}{p_{\min}}$, whereas both PMC and LDP values can grow arbitrarily large, even for secrets with finite alphabets (see Remark 3). As a result, meaningful comparisons between the three notions can only be made within the range $\varepsilon\in\left[0,\log\frac{1}{p_{\min}}\right]$.
Figure 1 presents the privacy-utility tradeoff curves for the randomized response mechanism across various priors and alphabet sizes, with utility measured via mutual information. In all cases, mechanisms adapted to PMC or PML achieve higher utility than those adapted to LDP, reflecting the advantage of context-aware guarantees over context-free ones. Between PMC and PML, the comparison depends on the prior and the privacy level. When $\varepsilon\leq\log\left(\frac{1-p_{\max}}{p_{\min}}\right)$, the PMC-adapted mechanism has a larger randomized response parameter ($\varepsilon_{r}$) than the PML-adapted one, and therefore yields higher utility. This effect is especially prominent for uniform priors over large alphabets. Conversely, when the alphabet is small and the prior is highly skewed, the PML-adapted mechanism offers higher utility.
(a) $n=3,P_{X}=(\frac{1}{3},\frac{1}{3},\frac{1}{3})$.
(b) $n=3,P_{X}=(0.8,0.1,0.1)$.
(c) $n=5,P_{X}=(0.6,0.1,0.1,0.1,0.1)$.
(d) $n=5,P_{X}=(0.8,0.05,0.05,0.05,0.05)$.
(e) $n=10,P_{X}$ uniform.
(f) $n=100,P_{X}$ uniform.
Figure 1: Privacy–utility tradeoff of the generalized randomized response mechanism. LDP-adapted, PMC-adapted, and PML-adapted mechanisms are compared under different priors, with utility measured by mutual information.
@@
Suppose $X$ is a finite random variable with alphabet $\mathcal{X}=[n]$. Recall from Section II-B that PML-extremal mechanisms constitute a class of optimal mechanisms maximizing sub-convex utility functions [25, Def. 5]. Let $0\leq\varepsilon_{u}<\log\,\frac{1}{1-p_{\mathrm{min}}}$. This range of values for $\varepsilon_{u}$ corresponds to the high-privacy regime. Extremal mechanisms in the high-privacy regime can be expressed as
@@
$\Lambda(X\to j)$ $=\log\;\max_{i\in[n]}\;\frac{P_{Y}(j)}{P_{Y\mid X=i}(j)}$
$=\log\,\frac{P_{X}(j)}{1-e^{\varepsilon_{u}}\Big(1-P_{X}(j)\Big)}.$
@@
Below, we evaluate the PMC from each data point to the released outcome. Due to the symmetry of the problem, the Laplace mechanism leaks the same amount of information about all data points, and for notational convenience, we calculate $\Lambda(X_{n}\to y)$ with $y\in\mathbb{R}$. [^2] While calculating $\Lambda(X_{n}\to y)$, the randomness of $X^{n-1}=(X_{1},\ldots,X_{n-1})$ contributes to the randomness of the mechanism $P_{Y\mid X_{n}}$. Specifically, we have $f_{Y\mid X_{n}}(y\mid x_{n})=\int f_{Y\mid X^{n}}(y\mid x_{1},\ldots,x_{n})\;P_{X^{n-1}}(dx_{1},\ldots,dx_{n-1}).$ Furthermore, a bound $\sup_{y}\Lambda(X_{n}\to y)\leq\varepsilon$ implies that the mechanism $P_{Y\mid X_{n}}$ satisfies $\varepsilon$-PMC under the prior $P_{X_{n}}$. Our analysis in this example complements a similar analysis performed with the Laplace mechanism and PML in [56, Prop. 4.5].
@@
First, we argue that to find the largest PMC across the $y$’s, it suffices to consider $y\geq d$ and $y\leq c$. To see why, note that in the numerator, using the fact that $\lvert a\rvert\leq\min\{-a,a\}$ for all $a\in\mathbb{R}$ we have
$\mathbb{E}_{X^{n}}\!\Bigg[\exp\!\Bigg(\!-\frac{\Big\lvert y-\frac{\sum_{i=1}^{n}X_{i}}{n}\Big\rvert}{b}\Bigg)\!\Bigg]\!\leq$
$\!\min\!\left\{\mathbb{E}\Bigg[\!\exp\!\Bigg(\!-\frac{y-\frac{\sum\limits_{i=1}^{n}X_{i}}{n}}{b}\!\Bigg)\!\Bigg],\mathbb{E}\Bigg[\!\exp\!\Bigg(\frac{y-\frac{\sum\limits_{i=1}^{n}X_{i}}{n}}{b}\!\Bigg)\!\Bigg]\right\},$
for all $y\in\mathbb{R}$. Moreover, in the denominator, the mapping
$y\mapsto\mathbb{E}_{X^{n-1}}\Bigg[\exp\Bigg(-\frac{\lvert y-\frac{x_{n}}{n}-\frac{\sum_{i=1}^{n-1}X_{i}}{n}\rvert}{b}\Bigg)\Bigg],$
is increasing in $(-\infty,c]$ and decreasing in $d,\infty)$ for all $x_{n}\in[c,d]$.
@@
$\mathbb{E}\!\left[\exp\!\left(-\frac{y-\frac{\sum_{i=1}^{n}X_{i}}{n}}{b}\right)\right]\!=\!\exp\!\left(-\frac{y}{b}\right)\mathbb{E}\!\left[\exp\!\left(\frac{\sum_{i=1}^{n}X_{i}}{nb}\right)\right]$
@@
$\mathbb{E}_{X^{n-1}}\left[\exp\left(-\frac{\displaystyle y-\frac{x_{n}}{n}-\frac{\sum_{i=1}^{n-1}X_{i}}{n}}{b}\right)\right]$
$=\exp\left(-\frac{y}{b}\right)\cdot\exp\left(\frac{x_{n}}{nb}\right)\;\mathbb{E}_{X^{n-1}}\left[\exp\left(\frac{\sum_{i=1}^{n-1}X_{i}}{nb}\right)\right]$
@@
$K(\frac{1}{nb})=K(-\frac{1}{nb})=\log\int_{\frac{c-d}{2}}^{\frac{d-c}{2}}\frac{1}{d-c}\exp(\frac{x}{nb})\,dx$
@@
$\log\;\mathbb{E}_{X}\left[\exp\left(\frac{2y(A+X)}{2\sigma^{2}}\right)\right]\leq\Lambda(X\to y)\\ \leq\log\;\mathbb{E}_{X}\left[\exp\left(\frac{2y(A+X)+A^{2}}{2\sigma^{2}}\right)\right],$
@@
$\frac{Ay}{\sigma^{2}}\leq\Lambda(X\to y)\leq\frac{A(A-4y)}{2\sigma^{2}}.$ (45)
@@
$\mathbb{P}\Big\{\Lambda(X\to Y)\geq\beta+$ $\frac{A^{2}}{2\sigma^{2}}\Big\}$ (47)
$\leq\mathbb{P}\left\{\frac{A(A+4\lvert Y\rvert)}{2\sigma^{2}}\geq\beta+\frac{A^{2}}{2\sigma^{2}}\right\}$ (48)
$=\mathbb{P}\left\{\lvert Y\rvert\geq\frac{\sigma^{2}\beta}{2A}\right\}$ (49)
@@ V PMC, PML, and Connections to Other Privacy Measures
In general, the upper and lower bounds on the information density are not independent quantities. These bounds influence each other so that imposing a lower bound on the information density may imply a certain upper bound, and vice versa. In the context of privacy, this naturally raises the question of whether a PMC guarantee inherently provides a certain PML guarantee or whether a PML guarantee implies a certain PMC guarantee.
In this section, we explore how PMC and PML guarantees interact with each other and with other privacy definitions, namely, ALIP, LIP, and LDP. First, in Section V-A, we assume that $X$ is a finite random variable and discuss how various privacy guarantees interrelate. We also highlight the implications of our results for designing mechanisms with optimal utility. Then, in Section V-B, we extend our analysis to secrets on standard Borel spaces.
@@ V-A Secrets with Finite Alphabets
Suppose $X\sim P_{X}$ is a finite random variable, and recall that $p_{\mathrm{min}}=\min_{x\in\mathcal{X}}P_{X}(x)$. We begin by presenting two results that illustrate how LDP and PMC relate. The first parallels Proposition 1 by showing how LDP constrains the information density through a bound on PMC. The second parallels Theorem 1 and establishes an equivalence between LDP and PMC over all full-support prior distributions. See Appendices E and F for proofs.
@@
The following result expresses LDP as a PMC constraint and provides a novel interpretation of LDP in terms of risk-averse adversaries.
Theorem 5.
Let $\mathcal{P}_{\mathcal{X}}$ denote the set of distributions with full support on $\mathcal{X}$. For any $\varepsilon\geq 0$, the mechanism $P_{Y\mid X}$ satisfies $\varepsilon$-LDP, if and only if it satisfies $\varepsilon$-PMC for all $P_{X}\in\mathcal{P}_{\mathcal{X}}$.
@@
Theorem 6 is proved in Appendix G.
@@
The condition $\varepsilon_{u}<\log\frac{1}{1-p_{\min}}$ restricts us to the high-privacy regime of PML and is necessary for $\varepsilon_{l}^{*}$ to be well-defined. This is because a mechanism $P_{Y\mid X}$ satisfying $\varepsilon_{u}$-PML with $\varepsilon_{u}\geq\log\frac{1}{1-p_{\min}}$ may assign $P_{Y\mid X=x}(y)=0$ for some $(x,y)$ pair with positive probability. Such a mechanism does not satisfy $\varepsilon_{l}$-PMC for any finite value of $\varepsilon_{l}$. This is also why an $\varepsilon_{u}$-PML guarantee yields useful LIP, ALIP and LDP guarantees only if $\varepsilon_{u}<\log\frac{1}{1-p_{\min}}$ (see also Corollary 3).
@@
Figure 2 illustrates the bounds presented in Theorem 6. Consider a pair $(\varepsilon_{u},\varepsilon_{l})\geq(0,0)$ with $\varepsilon_{u}<\log\frac{1}{1-p_{\mathrm{min}}}$. Theorem 6 states that if $\varepsilon_{l}>\varepsilon_{l}^{*}$, then $\varepsilon_{u}$-PML is a strictly stronger privacy guarantee compared to $\varepsilon_{l}$-PMC. Such pairs constitute the gray area above the blue curve in Figure 2(a). Similarly, if $\varepsilon_{u}>\varepsilon_{u}^{*}$, then $\varepsilon_{l}$-PMC is a strictly stronger privacy guarantee compared to $\varepsilon_{u}$-PML. These pairs correspond to the gray area below the orange curve in Figure 2(a). In contrast, the blue region in Figure 2(a) represents $(\varepsilon_{u},\varepsilon_{l})$ pairs that satisfy $\varepsilon_{u}^{*}\geq\varepsilon_{u}$ and $\varepsilon_{l}^{*}\geq\varepsilon_{l}$. In this region, neither the $\varepsilon_{u}$-PML nor the $\varepsilon_{l}$-PMC privacy guarantee is stronger than the other. Interestingly, when $X$ is an equiprobable Bernoulli random variable, the gap between the $\varepsilon_{u}$-PML and $\varepsilon_{l}$-PMC curves (i.e., the blue region) disappears, as illustrated in Figure 2(b). We formalize this observation in Corollary 4.
@@
Let us now discuss some of the implications of Theorem 6. The primary consequence of this theorem is that both PML and PMC privacy guarantees translate into ALIP, LIP, and LDP guarantees. The ALIP and LIP guarantees simply follow from their definitions, and the LDP guarantees follow from [71, Prop. 1], where it is shown that $(\varepsilon_{l},\varepsilon_{u})$-ALIP implies $(\varepsilon_{l}+\varepsilon_{u})$-LDP.
@@
    $(\varepsilon_{l}^{*},\varepsilon_{u})\text{-ALIP},\quad\max\{\varepsilon^{*}_{l},\varepsilon_{u}\}\text{-LIP},\quad(\varepsilon_{l}^{*}+\varepsilon_{u})\text{-LDP}.$
  2. (ii)
@@
    $(\varepsilon_{l},\varepsilon_{u}^{*})\text{-ALIP},\quad\max\{\varepsilon_{l},\varepsilon_{u}^{*}\}\text{-LIP},\quad(\varepsilon_{l}+\varepsilon_{u}^{*})\text{-LDP}.$